### v1

# Large Language Model-Enabled Multi-Agent Manufacturing Systems

Abstract
Traditional manufacturing faces challenges adapting to dynamic environments and quickly responding to manufacturing changes. The use of multi-agent systems has improved adaptability and coordination but requires further advancements in rapid human instruction comprehension, operational adaptability, and coordination through natural language integration. Large language models like GPT-3.5 and GPT-4 enhance multi-agent manufacturing systems by enabling agents to communicate in natural language and interpret human instructions for decision-making. This research introduces a novel framework where large language models enhance the capabilities of agents in manufacturing, making them more adaptable, and capable of processing context-specific instructions. A case study demonstrates the practical application of this framework, showing how agents can effectively communicate, understand tasks, and execute manufacturing processes, including precise G-code allocation among agents. The findings highlight the importance of continuous large language model integration into multi-agent manufacturing systems and the development of sophisticated agent communication protocols for a more flexible manufacturing system.
I Introduction
The escalating demand for customized products and the increasing complexities of production processes are driving a transformation in the manufacturing sector. This paradigm shift has required manufacturers to quickly adjust to evolving product specifications and adeptly address unforeseen operational changes. In this context, manufacturers are required to be not only agile and responsive but also to maintain high-quality products and cost efficiency in a dynamically changing environment.
One significant approach for enhancing the adaptability and quick response of manufacturing systems is the use of multi-agent control [1, 2]. A Multi-Agent System (MAS) is composed of several autonomous agents for decision-making, communication, and coordination [3]. Many multi-agent architectures have been developed for controlling manufacturing systems [1, 4, 5, 6]. By enabling decentralized decision-making and enhancing coordination among various agents, MAS have contributed significantly to handling complexities in the manufacturing system [7].
However, there are still challenges in integrating MAS frameworks in the manufacturing context. While agents in MAS can automate processes and make decisions from available data, their ability to adapt to new formats or specifications is limited, leading to increased downtime. Thus, a significant challenge includes the limited ability of agents to dynamically adapt to new information including product specifications (RCh1). Another challenge lies in processing unstructured textual data. Textual data is crucial because it is more expressive than numerical data, offering vast knowledge that is currently underutilized in manufacturing, where structured data fails to include varied influencing factors [8]. The capability of natural language processing to formalize unstructured textual data can enable downtime reduction and process optimization [8]. Therefore, advanced natural language processing is essential for interpreting and implementing textual data from human operators or interconnected systems (RCh2).
The advent of Large Language Models (LLMs) presents a novel solution to these challenges. LLMs, such as OpenAI’s GPT-3 [9] and GPT-4 [10], provides opportunities to bring more flexibility and adaptability into multi-agent manufacturing systems. With Machine to Machine (M2M) protocols in the manufacturing industry, such as Message Queuing Telemetry Transport (MQTT) [11] and MTConnect [12] that uses predefined structures, incorporating LLMs offer benefits. LLMs can interpret, generate, and act on natural language instructions, facilitating complex decision-making processes and adaptive responses to changing manufacturing conditions [13]. The integration of natural language capabilities may also benefit operators, engineers, and other system users who have limited experience with the specific process or system [14, 15]. For manufacturing systems, natural language capabilities can allow new users to quickly and efficiently adapt to managing and optimizing processes and resources [16]. The main contributions of this paper include: Leveraging LLM to allocate and execute manufacturing tasks based on user instructions, to enhance task efficiency (Cn1), enabling LLM agents to interpret and adapt to evolving specifications ensuring the production process is correctly executed with the appropriate resources (Cn2), and applying advanced natural language processing capabilities to enable agents to effectively comprehend context-specific instructions from users and other agents (Cn3).
The rest of the manuscript is organized as follows. Section II provides background regarding MAS, LLMs in MAS, and LLMs in manufacturing. Section III details the architecture and functioning of the proposed framework. Section IV demonstrates a case study to illustrate the applications and effectiveness of our framework. Section V presents challenges and limitations encountered while integrating LLM. Section VI summarizes the paper and discusses future work.
II Background
II-A Multi-Agent Systems in Manufacturing
Various MAS architectures have been developed to enhance agent communication, significantly contributing to the increased flexibility of manufacturing systems. For instance, the Axiomatic Design of a Multi-Agent Reconfigurable Mechatronic System (ADMARMS) framework is an early MAS architecture designed for reconfigurable manufacturing, incorporating axiomatic design and quantitative methodologies to facilitate design-based reconfigurability. Studies by Wang et al. [17] and Kim et al. [18] have expanded the scope of MAS by integrating these systems with big data analytics and reinforcement learning. Wang et al. focused on improving manufacturing flexibility through agent negotiation for deadlock prevention, while Kim et al. aimed at facilitating autonomous decision-making, particularly for personalized production systems. Additionally, the works of Kovalenko et al. [19] and Bi et al. [6] have concentrated on dynamic resource and task allocation within MAS. Kovalenko et al. introduced a framework for dynamic resource task negotiation, utilizing a network-based coordination strategy that enables Product Agents (PAs) to effectively explore and understand resource capabilities. Similarly, Bi et al. developed a model-based Resource Agent (RA) architecture to address disruptions in manufacturing systems by improving coordination and decision-making capabilities.
While these MAS architectures were developed to enhance agent communication for increased flexibility, they focus on utilizing predefined mathematical models and algorithms for task allocation and agent communication. Furthermore, there has been limited exploration in using natural language interpretation from human operators or other agents to enhance the responsiveness of manufacturing process control. By leveraging LLMs, agents can achieve higher operational flexibility and efficiency in addressing these gaps.
II-B Large Language Models in Multi-Agent Systems
The emergence of LLMs has sparked significant interest in the development of multiple interacting autonomous agents powered by the LLM technology [20]. The agents enabled by LLMs demonstrate enhanced problem-solving capabilities through multi-agent discussions and diverse applications, from sociological studies to complex task solutions. For instance, Park et al. [21] showcase a simulation with 25 generative agents in a virtual town, designed to exhibit realistic human behavior and facilitate studying social understanding through memory retrieval. Another work by Wu et al. [22] introduces an open-source framework called AutoGen that facilitates customizable agent interactions and conversation programming. ChatEval [23] incorporates debater agents in an LLM-powered MAS for text evaluation, achieving enhanced accuracy and depth over traditional methods and single agent setups. LLMs have also been effectively integrated into specialized domains like software engineering. MetaGPT [24], a meta-programming framework, enhances LLM-based multi-agent collaboration using an assembly line approach to assign different roles to each agent for software engineering tasks, outperforming previous chat-based systems in complex problems. Similarly, ChatDev  [25], a virtual chat-powered framework that applies LLMs to automate and streamline the entire software development process, including efficient collaboration across designing, coding, testing, and documenting phases.
These works demonstrate advanced multi-agent interactions, realistic behavior simulation, and efficient tasks, which can be applied in manufacturing to enhance adaptability, coordination, and advanced natural language processing capabilities. Building on the insights from prior studies, we delve into the integration of LLMs in a multi-agent framework, aiming to address the unique requirements and benefits in the manufacturing sector.
II-C Large Language Models in Manufacturing
With the advancement of LLM technology, recent studies have also been increasingly focusing on exploring both the opportunities and challenges associated with implementing LLMs within the manufacturing domain. Makatura et al. [13] explores the application of LLMs in design and manufacturing, including CAD/CAM. Their approach investigates the potential role of LLMs in a variety of tasks, including design specification and manufacturing instruction, and highlights the capabilities and limitations. The analysis of GPT-4 shows strengths in knowledge and iteration in design and manufacturing but faces challenges in reasoning and accuracy, with potential solutions like specialized languages and improved context management [13]. Furthermore, recent research on LLM in manufacturing delves into language models for specific manufacturing applications. Jignasu et al. [26] presents a comprehensive evaluation of LLMs for understanding, debugging, and manipulating G-code in 3D printing, testing their capabilities in error detection and correction, geometric transformations, and discussing their potential and limitations in G-code comprehension. Similar work has been done by Badini et al. [27], which explores using ChatGPT to optimize the G-code generation in additive manufacturing, enhancing efficiency and quality in manufacturing processes. These studies prove that LLM can interpret G-code, essential for tasks like debugging, manipulating, and comprehending the code used in manufacturing processes. Studies by Fakih et al. [28] and Koziolek et al. [29] explore the use of LLMs for enhancing programming in industrial control systems, specifically focusing on Programmable Logic Controllers (PLCs). Fakih et al. developed LLM4PLC, a framework that iteratively refines LLM-generated code through user feedback and verification tools to improve code generation quality. Koziolek et al. conducted an exploratory study to assess GPT-4’s capability to generate control logic from natural language prompts.
Several studies have investigated the integration of LLMs into manufacturing, including areas such as CAD/CAM, G-code, and PLCs, aiming to enhance efficiency in manufacturing processes. However, these studies have not focused on using LLMs with MAS for autonomous task allocation, adaptability to evolving specifications, and interpretation of context-specific instructions. The framework described in Section III focuses on developing autonomous agents that use instructions and specifications in context-specific natural language to dynamically adjust to new product information and coordinate during runtime to complete manufacturing tasks. Our proposed framework can be identified as an initial step toward enhancing the CAD/CAM workflow for automatically adapting CAM processes based on changes in CAD specifications, ensuring seamless manufacturing operations.
III Framework
Figure 1: Framework for LLM-Enabled Multi-Agent Manufacturing Systems
The novel architecture presented in Figure 1 is designed based on the architecture of a MAS that includes both PAs and RAs. A PA is an agent responsible for decision-making related to a product, while an RA serves as a controller for a specific resource, such as a robot or machine [30].
Figure 1 provides a general overview of an LLM-enabled multi-agent manufacturing system framework, where a user initializes LLM agents with instructions and capabilities. PAs are assigned specific processes including G-code specifications, while possessing the ability to communicate and access historical data for informed decision-making. RAs are responsible for the process execution of tasks, informed by instructions and tool specifications provided by the user. LLM is applied for processing and responding to natural language inputs from both PAs and RAs. For evaluation, the chosen LLM is GPT-4, the latest and most advanced version of GPT developed by OpenAI [10].
This framework facilitates adaptability, enables real-time coordination, and provides the ability to respond to the demands of a complex manufacturing environment with natural language processing capabilities.
III-A Agent Initialization
Algorithm 1 Agent Creation
1:procedure CreateAgents($Agent\_list$, $Agent\_functions$)
2:  $Agents\leftarrow[]$
3:  for $agent\_initialization$ in $Agent\_list$ do
4:    $configuration\leftarrow$ LoadJSON($agent\_initialization$)
5:    for $name$, $agent\_data$ in $configuration$ do
6:     $function\_list\leftarrow[]$
7:     if ’functions’ in $agent\_data$ then
8:       for $function\_name$ in $agent\_data$[’functions’] do
9:         $function\_reference\leftarrow$ GetFunction($Agent\_functions$,
10:            $function\_name$)
11:         if $function\_reference\neq$ None then
12:          $function\_list$.append($function\_reference$)
13:     $agent\leftarrow$ NewAgent($name$, $agent\_data$, $function\_list$)
14:     $Agents$.append($agent$)
15:  return $Agents$
The agent initialization process aims to configure a cooperative MAS for operation by defining roles, capabilities, and communication protocols for each agent. This setup ensures agents operate effectively and collaborate towards system goals. The following initial prompt is given to all agents in the system:
"You are a helpful agent in a cooperative Multi-Agent System. If you are asked for a service you can provide you should help. If necessary, you may ask the other agent for clarifying information. You may communicate with your peers to achieve your goals. If you do not know the answer do not make things up. Only use the functions you have been provided with. However, you may call these functions recursively."
These instructions ensure that each agent operates within the boundaries of its capabilities and knowledge, engages in effective communication, and contributes to the overall objectives of the MAS.
Following the initial prompts common to all agents in the system, the initialization process involves various parameters, assigning each agent to its specific role. These parameters were selected to achieve the desired behaviors of adaptability, real-time response, autonomous decision-making, and efficient coordination within the MAS. The process includes the following key parameters essential for the agents:
  • Functions: Capabilities within the manufacturing process, that range from retrieving manufacturing history to executing manufacturing tasks.
  • Annotation: Agent’s role and responsibilities, aiding the selection of the appropriate agent among other LLM agents.
  • Instructions: User instructions offer a sequence of operations, techniques, and standards for agent activities.
  • Inbox: The agent is initialized with an inbox to receive and manage instructions, updates, and feedback from various agents.
For PAs, an additional parameter is introduced:
  • Specification File: Detailed manufacturing specifications, such as G-code files for CNC operations, for precise manufacturing execution.
RAs are further configured with:
  • Configurations: This parameter outlines the operational settings specific to RAs, such as process times, defect rates, and breakdown probabilities.
The integration of these parameters into LLM-enabled MAS for manufacturing extends the structured and standardized approaches (e.g. FIPA [31] and JADE [32]) to incorporate natural language processing capabilities for more flexible and human-like agent interactions. Integrating user-defined instructions and parameters in agent initialization can be further expanded through user studies to ensure applicability to real scenarios and industries.
Based on the initialization information, the Agent Creation function, cp. Algorithm 1, enables the dynamic generation of agents. The Agent Creation function provides advanced LLM agents with different capabilities in manufacturing systems, improving operational efficiency and adaptability to changing requirements. The algorithm iterates over each entry in the Agent_list, which contains initialization information. For every entry, configuration data configuration for the agent is loaded from a JSON, containing details such as the agent’s name, role, capabilities, and any specific functions the agent can perform. For each agent specified in the configuration data, an empty list function_list is initially prepared. If the ‘functions’ key is present in the agent’s data, the algorithm iterates through each function name listed under this key. For each function name, a reference function_reference is obtained by searching within a provided set of agent functions Agent_functions. If the function is found, its reference is appended to function_list. A new agent instance agent is created with the specified name, agent data, and the list of function references. This agent is then added to the Agents list. This algorithm facilitates the flexible creation of agents allowing for dynamic adaptation to various tasks and scenarios.
Algorithm 2 Chat Interaction
1:procedure chat($prompt$, $model$)
2:  $with\_functions\leftarrow$ len(function_info) $>0$
3:  $msgs\leftarrow[]$
4:  if instructions then
5:    msgs.append({"role": "system", "content": instructions})
6:  msgs.append({"role": "user", "content": $prompt$})
7:  $response\leftarrow$_get_response($msgs$, $model\_$, $with\_functions$)
8:  msgs.append($response$.message)
9:  $function\leftarrow response$.message.function_call.name
10:  $args\leftarrow$ json.loads($response$.message.function_call.arguments)
11:  $function\_response\leftarrow$ executables[$function$](**$args$)
12:  return $function\_response$, $msgs$
III-B Agent Capabilities and LLM Integration
Through the integration of LLM, agents gain the capability to interpret complex instructions and specifications, facilitating a flexible approach to manufacturing operations. The framework employs LLM to enable agents to decode user instructions for manufacturing, integrating a communication mechanism through function information and execution capabilities with the OpenAI Application Programming Interface (API) feature called function calling [33]. This allows for the automatic execution of user-defined functions, including specific manufacturing operations (e.g., milling, drilling) through descriptions that include operation parameters and required inputs like product name, machine name, process name, and product specifications (e.g. G-code).
The ‘chat’ function, cp. Algorithm 2, facilitates dynamic interaction among agents, based on available function information and executing function calls as defined by the manufacturing process requirements, utilizing the appropriate LLM model. The outcome, function_response, includes the details like operation type, product, and machine names, along with a message, msgs, to facilitate the subsequent steps in the manufacturing process. This approach enables context-aware execution of manufacturing tasks, enhancing the adaptability and efficiency of the manufacturing process.
III-C Operational Workflow
Figure 2: Operational Workflow of the Framework Using LLM for Task Execution and Coordination
The workflow, as depicted in Figure 2, shows how user instructions in textual data can be converted into actionable manufacturing operations. This approach enables the translation of high-level directives of users into machine-executable actions, enhancing the responsiveness of the manufacturing workflow to changing conditions.
The workflow begins with the user initializing PAs and RAs within the manufacturing system. Users describe product goals, sequences of operations, and specifications. This information is then processed by the LLM, which generates responses to guide the process. The PA determines the suitable RA for the next task, prompting it to carry out the required operation. The selected RA constructs the necessary arguments to request the corresponding function call, leading to the execution of the specified process. Upon successful operation, the LLM formulates an appropriate response, informing the PA of the process status. This feedback updates the LLM-integrated PA with the current process status, determining the completion of the task. If additional tasks remain, the cycle repeats; otherwise, the user is notified with the completion of all processes.
IV Case Study
In this case study, we examine the proposed framework and demonstrate how the LLM is integrated into a multi-agent framework to execute the manufacturing process.
IV-A Case Study Setup
Figure 3: Case Study Setup
A simulated manufacturing environment is equipped with three distinct machines, each specialized for milling, drilling, and threading tasks, as highlighted in Figure 3. To evaluate the system’s adaptability to new product information, we structured the case study into two distinct processes: a 2-step process and a more comprehensive 4-step process. In the 2-step process, a prototype product, referred to as ‘product-1’, is subjected to contouring and drilling operations. The 4-step process extends this sequence by incorporating counterboring and threading operations, testing the system’s capability to handle extended manufacturing workflows. The G-code specifications for these processes are adapted from [34], as detailed in Figure 4. Notably, explicit labels for G-code specifications (e.g. Step1-Contouring, Step2-Drilling) are not provided to the LLM agent. Leveraging LLM’s capability to interpret and assign appropriate G-code based on tool specifications for each resource agent demonstrates the system’s potential to efficiently manage G-code assignments, demonstrating the advanced comprehension capabilities of LLMs in a simulated manufacturing context.
Each product has an LLM-enabled PA which oversees high-level decision-making for its associated product. For instance, ‘product-1’ is instructed for a 2-step process with the following instructions as:
"Make sure to get contouring to shape the external profile of the product. Then, get a drilling operation to create precise holes. This operation can precede or follow the contouring. Proceed to Exit Buffer completed. Your product name is ’product-1’. The G-code specifications are as follows: %358 N1 N10 G90 G71 G80 G40..."
Likewise, each LLM-enabled RA is responsible for operational decisions related to its specific machine. ‘Milling1’ is tasked with both contouring and counterboring to demonstrate how LLM can adaptively assign operations to the appropriate machine. ‘Drilling1’ focuses on drilling, while ‘Threading1’ manages threading tasks. Each machine is assigned specific tool numbers that are used in G-code specifications to ensure that PAs provide only the relevant G-code instructions to each RA. For example, the ‘Milling1’ machine receives the following operational prompt:
"Perform milling operations as required, including contouring the external shape and counterboring for holes. Your machine name is Milling1. The tool number for contouring T4. The tool number for counterboring is T6."
This experiment investigates the LLM’s capability to identify and allocate correct manufacturing task and their corresponding G-code specifications. It aims to enhance the adaptability by showcasing natural language communication between PA and RA, simulating human-like interaction for seamless task execution within the manufacturing environment.
Figure 4: Case Study G-code Product Specification
IV-B Case Study Simulation and Results Analysis
Figure 5: Agent Communication Result for 2-step Process
Following the setup of our case study, simulations were conducted to evaluate the system performance, as shown in one example, depicted in Figure 5. In this example, the user initiates the workflow by instructing the ‘Product-1’ agent and all other RAs, using  Algorithm 1 from Section III-A. Following the agent initialization, the system operates according to the steps detailed in the operational workflow, as depicted in Section III-C. First, the ‘Product-1’ agent communicates the need for contouring to the ’Milling1’ agent, providing only the contouring section of G-code specifications. After the ‘Milling1’ agent completes the task, it informs the ‘Product-1’ agent, indicating it is ready for the next operation. The process continues with the ‘Product-1’ agent assigning the drilling operation to the ‘Drilling1’ agent, which updates the ‘Product-1’ agent after completion. When all operations are completed, the ‘Product-1’ agent updates the user, indicating that all manufacturing processes for ‘Product-1’ are completed and are now ready to move to the exit buffer. All agent communications are facilitated by the chat interaction protocol defined in Algorithm 2, as outlined in Section III-B.
This section details the outcomes from the simulations, highlighting the communications among agents, their understanding and execution of tasks, the system’s accuracy, and the insights gained from these case studies.
IV-B1 Communication and Comprehension
Our case study focused on evaluating the agents’ ability to communicate and comprehend specific tasks within the manufacturing process. Utilizing natural language processing, agents interacted in both the 2-step and 4-step processes, showcasing their capability for nuanced communication. In the 2-step process, the communication between the ’Product-1’ agent and the ‘Milling1’ and ‘Drilling1’ agents showed the system’s efficiency in handling straightforward task sequences and distributing G-code specifications in human-like communication. The 4-step process further explored the agents’ capability in managing and understanding complex sequences, requiring accurate G-code allocation where ‘Milling1’ was tasked with both contouring and counterboring. In this scenario, the ‘Product-1’ agent coordinated the sequence of contouring, drilling, counterboring, and threading, showing the agents’ ability to interpret and implement longer instructions.
IV-B2 Accuracy and Performance
The data presented in Table I reveals that GPT-4 achieved success rates of 100% for 2-step and 86% for 4-step manufacturing processes across 50 trials for each. This shows a decline in performance as task complexity increases, with the 4-step process experiencing a 14% overall error rate. The detailed error breakdown in  Table II reveals that within the overall 14% error rate for the 4-step process, incorrect function calls and process inexecution have 43% of the errors, while inaccurate G-code allocation accounts for 14% of the total errors. It is observed that the increase in prompt length correlates with a rise in errors, highlighting the need for refined model training and improved prompt engineering. Addressing these issues aims to enhance task execution capabilities and effective integration of LLMs into MASs for manufacturing, improving system adaptability and operational efficiency even in unexpected scenarios.
IV-B3 Insights from Case Studies
The case study provided insights into the capabilities of LLM-enabled MASs in manufacturing. The current framework incorporates LLM to transition towards a system capable of understanding the manufacturing tasks (Cn1), interpreting and adapting to new product specifications (Cn2), and agent communication via natural language instructions (Cn3). This shift enhances system flexibility and expands its ability to manage various tasks without relying on predetermined models or rules. However, accuracy issues with GPT-4 highlight the need for model improvement and adaptation to manufacturing needs. Difficulties with incorrect argument passing during function calls point to the necessity for better prompt engineering, error management, and validation within the system. These results guide future research toward enhancing LLM integration in manufacturing, developing advanced agent communication protocols, and increasing system robustness against errors.
TABLE I: Success Rates for Case Studies with GPT-4
Metric 2-Step 4-Step
Success Rate 1 .86
Overall Error Rate 0 .14
- P1. Contouring Failure Rate 0 0
- P2. Drilling Failure Rate 0 .06
- P3. Counterboring Failure Rate N/A .02
- P4. Threading Failure Rate N/A .06
TABLE II: Detailed Error Analysis with GPT-4
Error Type 2-Step 4-Step
Incorrect Function Calls 0 .43
Inaccurate G-code Allocation 0 .14
Process Inexecution 0 .43
V Challenges and Limitations
V-A Performance
A significant challenge within the use of LLMs in manufacturing processes is achieving the level of performance required for high-quality production. Manufacturing processes demand precision to ensure product consistency, safety, and reliability. However, LLMs, while effective in processing natural language, occasionally shows limitations in delivering the precision necessary for these applications, as indicated in Section IV-B2. This issue raises concerns about the feasibility of relying on LLMs for tasks where errors can lead to other consequences, such as machine breakdown and low product quality.
V-B Scalability
Scalability is also a challenge when integrating LLMs into manufacturing systems. As the volume of data and the complexity of tasks grow, there is a risk that LLMs may not maintain their initial levels of accuracy and efficiency. This limitation is particularly noticeable in situations where extensive historical data and detailed prompts are essential for task execution, as exemplified in the 4-step process in Section IV. The potential decrease in accuracy with larger datasets poses a challenge to scaling LLM-enabled multi-agent control across more complex manufacturing operations.
V-C Reliability
The integration of LLMs into manufacturing settings highlights the importance of reliability. LLMs generating hallucinated responses or making false decisions raises significant safety concerns, as shown in Section IV. The LLM in the case study allocated inaccurate G-code or executed the incorrect function calls. Moreover, the reasoning provided by LLMs may require validation in practical manufacturing scenarios even if it seems logical. Therefore, it is essential to implement robust verification mechanisms, as well as incorporate real-time observability and monitoring capabilities. This approach ensures LLMs’ outputs and reasoning are consistent and allow immediate adjustments with manufacturing standards and safety protocols. For instance, the study by Joublin et al. [35] presents a system architecture for robotics that utilizes LLMs for planning tasks, focusing on a strategy for replanning in response to errors via feedback. This method enables robots to adjust to failures and changes, enhancing reliability. Thus, feedback loops in LLM systems ensure safety and reliability in complex manufacturing environments.
VI Conclusion and Future Work
In our research, we explored the integration of LLM within MAS to enhance manufacturing processes. Our research demonstrates the effect of LLM in manufacturing, where they serve to dynamically match tasks with optimal process parameters, improving task allocation efficiency and operational adaptability (Cn1). Moreover, LLM enables agents to adapt to evolving product specifications (e.g. G-code), meeting the manufacturing requirements (Cn2). Additionally, through advanced natural language processing, LLM enables agents to effectively understand and execute complex, context-specific instructions from humans and other agents enhancing the agent communication and collaboration within the manufacturing process (Cn3). However, challenges have been identified, such as task execution performance, scalability, and reliability, highlighting the need for a verification process and monitoring capabilities when making decisions for manufacturing operations. Furthermore, the diversity in user instructions based on different skill levels and manufacturing requirements affects task interpretation and execution, which requires a strategy to dynamically adapt to these changes when integrating LLM to ensure reliability.
Future work will focus on developing communication protocols incorporating LLM for agents that can accommodate a wide variety of manufacturing instructions and requirements, influenced by varying proficiency levels of human operators. This approach seeks to showcase the capability of agents to learn and adapt from historical data, enhancing their responsiveness to rapid changes in modern manufacturing industries. Additionally, efforts will be directed towards enhancing the accuracy and scalability of LLM. While this paper initially focused on widely used manufacturing processes (e.g. drilling, milling), future investigations will extend to include less commonly used manufacturing processes with larger datasets to evaluate the effectiveness of our framework. The goal is to assess if integrating additional processes results in less effective outcomes or if they can also be improved by applying these methodologies, ensuring efficient processing of larger datasets and diverse manufacturing processes without impacting performance.
Acknowledgment
We thank Felix Ocker for his valuable comments and suggestions.

### v2

@@
 This paradigm shift has required manufacturers to quickly adjust to evolving product specifications and adeptly address unforeseen operational changes.
-In this context, manufacturers are required to be not only agile and responsive but also to maintain high-quality products and cost efficiency in a dynamically changing environment.
-One significant approach for enhancing the adaptability and quick response of manufacturing systems is the use of multi-agent control [1, 2].
-A Multi-Agent System (MAS) is composed of several autonomous agents for decision-making, communication, and coordination [3].
-Many multi-agent architectures have been developed for controlling manufacturing systems [1, 4, 5, 6].
+In this context, manufacturers are required to not only maintain high-quality products and cost efficiency but also be agile and responsive to a dynamically changing environment.
+One significant approach for enhancing adaptability and quick response of manufacturing systems is using multi-agent control [1].
+A Multi-Agent System (MAS) includes several autonomous agents for decision-making, communication, and coordination [2].
+Many multi-agent architectures have been developed for manufacturing systems [3, 4].
 By enabling decentralized decision-making and enhancing coordination among various agents, MAS have contributed significantly to handling complexities in the manufacturing system [5].
-However, there are still challenges in integrating MAS frameworks in the manufacturing context.
-While agents in MAS can automate processes and make decisions from available data, their ability to adapt to new formats or specifications is limited, leading to increased downtime.
-Thus, a significant challenge includes the limited ability of agents to dynamically adapt to new information including product specifications (RCh1).
+However, there are still challenges in MAS frameworks in manufacturing.
+While agents in MAS can automate processes and make decisions, their ability to adapt to new formats or specifications is limited, leading to increased downtime.
+Thus, a challenge includes the limited ability of agents to dynamically adapt to new information including product specifications (RCh1).
 Another challenge lies in processing unstructured textual data.
 Textual data is crucial because it is more expressive than numerical data, offering vast knowledge that is currently underutilized in manufacturing, where structured data fails to include varied influencing factors [6].
-The capability of natural language processing to formalize unstructured textual data can enable downtime reduction and process optimization [8].
-Therefore, advanced natural language processing is essential for interpreting and implementing textual data from human operators or interconnected systems (RCh2).
+Natural language processing capabilities to formalize unstructured textual data can reduce downtime and optimize processes [6].
+Therefore, natural language processing is essential for interpreting and implementing textual data from human operators or interconnected systems (RCh2).
 The advent of Large Language Models (LLMs) presents a novel solution to these challenges.
-LLMs, such as OpenAI’s GPT-3 [9] and GPT-4 [10], provides opportunities to bring more flexibility and adaptability into multi-agent manufacturing systems.
-With Machine to Machine (M2M) protocols in the manufacturing industry, such as Message Queuing Telemetry Transport (MQTT) [11] and MTConnect [12] that uses predefined structures, incorporating LLMs offer benefits.
+LLMs, such as OpenAI’s GPT-3 [7] and GPT-4 [8], provides opportunities to bring more flexibility and adaptability.
+With Machine to Machine (M2M) protocols, such as Message Queuing Telemetry Transport (MQTT) [9] and MTConnect [10] that uses predefined structures, incorporating LLMs offer benefits.
 LLMs can interpret, generate, and act on natural language instructions, facilitating complex decision-making processes and adaptive responses to changing manufacturing conditions [11].
-The integration of natural language capabilities may also benefit operators, engineers, and other system users who have limited experience with the specific process or system [14, 15].
-For manufacturing systems, natural language capabilities can allow new users to quickly and efficiently adapt to managing and optimizing processes and resources [16].
-The main contributions of this paper include: Leveraging LLM to allocate and execute manufacturing tasks based on user instructions, to enhance task efficiency (Cn1), enabling LLM agents to interpret and adapt to evolving specifications ensuring the production process is correctly executed with the appropriate resources (Cn2), and applying advanced natural language processing capabilities to enable agents to effectively comprehend context-specific instructions from users and other agents (Cn3).
+The integration of natural language processing capabilities may also benefit operators and other system users who have limited experience with the specific process or system [12].
+Natural language processing capabilities can allow new users to quickly adapt to managing and optimizing processes and resources [13].
+The main contributions of this paper include: Leveraging LLM to allocate and execute manufacturing tasks based on user instructions to enhance task efficiency (Cn1), enabling LLM agents to adapt to evolving specifications ensuring the production process is executed with the appropriate resources (Cn2), and applying natural language processing capabilities to enable agents to comprehend context-specific instructions from users and other agents (Cn3).
 The rest of the manuscript is organized as follows.
 Section II provides background regarding MAS, LLMs in MAS, and LLMs in manufacturing.
-Section III details the architecture and functioning of the proposed framework.
+Section III details the functioning of the proposed framework.
 Section IV demonstrates a case study to illustrate the applications and effectiveness of our framework.
-Section V presents challenges and limitations encountered while integrating LLM.
+Section V presents challenges and limitations while integrating LLM.
 Section VI summarizes the paper and discusses future work.
@@
 II-A Multi-Agent Systems in Manufacturing
-Various MAS architectures have been developed to enhance agent communication, significantly contributing to the increased flexibility of manufacturing systems.
-For instance, the Axiomatic Design of a Multi-Agent Reconfigurable Mechatronic System (ADMARMS) framework is an early MAS architecture designed for reconfigurable manufacturing, incorporating axiomatic design and quantitative methodologies to facilitate design-based reconfigurability.
+Various MAS architectures have been developed to enhance agent communication, significantly contributing to the increased manufacturing system flexibility.
+For instance, the Axiomatic Design of a Multi-Agent Reconfigurable Mechatronic System (ADMARMS) framework facilitates design-based reconfigurability [14].
 Studies by Wang et al.
 [15] and Kim et al.
-[18] have expanded the scope of MAS by integrating these systems with big data analytics and reinforcement learning.
-Wang et al.
-focused on improving manufacturing flexibility through agent negotiation for deadlock prevention, while Kim et al.
-aimed at facilitating autonomous decision-making, particularly for personalized production systems.
-Additionally, the works of Kovalenko et al.
+[16] integrated MAS with big data analytics and reinforcement learning.
+Additionally, Kovalenko et al.
 [17] and Bi et al.
-[6] have concentrated on dynamic resource and task allocation within MAS.
-Kovalenko et al.
-introduced a framework for dynamic resource task negotiation, utilizing a network-based coordination strategy that enables Product Agents (PAs) to effectively explore and understand resource capabilities.
-Similarly, Bi et al.
-developed a model-based Resource Agent (RA) architecture to address disruptions in manufacturing systems by improving coordination and decision-making capabilities.
-While these MAS architectures were developed to enhance agent communication for increased flexibility, they focus on utilizing predefined mathematical models and algorithms for task allocation and agent communication.
+[4] focused on dynamic resource and task allocation within MAS, with frameworks for resource task negotiation and a model-based resource agent architecture to address manufacturing disruptions.
+While these MAS architectures were developed to enhance agent communication for increased flexibility, they focus on utilizing predefined algorithms for task allocation and agent communication.
 Furthermore, there has been limited exploration in using natural language interpretation from human operators or other agents to enhance the responsiveness of manufacturing process control.
-By leveraging LLMs, agents can achieve higher operational flexibility and efficiency in addressing these gaps.
+By leveraging LLMs, agents can improve operational flexibility and efficiency in addressing these gaps.
 II-B Large Language Models in Multi-Agent Systems
-The emergence of LLMs has sparked significant interest in the development of multiple interacting autonomous agents powered by the LLM technology [20].
-The agents enabled by LLMs demonstrate enhanced problem-solving capabilities through multi-agent discussions and diverse applications, from sociological studies to complex task solutions.
+The emergence of LLMs has sparked interest in developing interacting autonomous agents powered by the LLM technology [18].
+LLM-enabled agents demonstrate enhanced problem-solving capabilities through multi-agent discussions and diverse applications.
 For instance, Park et al.
-[21] showcase a simulation with 25 generative agents in a virtual town, designed to exhibit realistic human behavior and facilitate studying social understanding through memory retrieval.
+[19] showcase a simulation with 25 generative agents in a virtual town, facilitating social understanding studies.
 Another work by Wu et al.
 [20] introduces an open-source framework called AutoGen that facilitates customizable agent interactions and conversation programming.
-ChatEval [23] incorporates debater agents in an LLM-powered MAS for text evaluation, achieving enhanced accuracy and depth over traditional methods and single agent setups.
-LLMs have also been effectively integrated into specialized domains like software engineering.
-MetaGPT [24], a meta-programming framework, enhances LLM-based multi-agent collaboration using an assembly line approach to assign different roles to each agent for software engineering tasks, outperforming previous chat-based systems in complex problems.
-Similarly, ChatDev  [25], a virtual chat-powered framework that applies LLMs to automate and streamline the entire software development process, including efficient collaboration across designing, coding, testing, and documenting phases.
-These works demonstrate advanced multi-agent interactions, realistic behavior simulation, and efficient tasks, which can be applied in manufacturing to enhance adaptability, coordination, and advanced natural language processing capabilities.
-Building on the insights from prior studies, we delve into the integration of LLMs in a multi-agent framework, aiming to address the unique requirements and benefits in the manufacturing sector.
+ChatEval [21] uses debater agents in an LLM-powered MAS for text evaluation, achieving enhanced accuracy and depth over traditional methods and single agent setups.
+MetaGPT [22], a meta-programming framework, enhances multi-agent collaboration using an assembly line approach for software engineering tasks.
+These works demonstrate advanced multi-agent interactions, realistic behavior simulation, and task efficiency, which can enhance adaptability, coordination, and advanced natural language processing capabilities in manufacturing.
+Building on these insights, we explore integrating LLMs in a multi-agent framework to address the unique requirements and benefits in the manufacturing sector.
 II-C Large Language Models in Manufacturing
-With the advancement of LLM technology, recent studies have also been increasingly focusing on exploring both the opportunities and challenges associated with implementing LLMs within the manufacturing domain.
+With the advancement of LLM technology, recent studies have been increasingly exploring both the opportunities and challenges of implementing LLMs within the manufacturing domain.
 Makatura et al.
@@
 Their approach investigates the potential role of LLMs in a variety of tasks, including design specification and manufacturing instruction, and highlights the capabilities and limitations.
-The analysis of GPT-4 shows strengths in knowledge and iteration in design and manufacturing but faces challenges in reasoning and accuracy, with potential solutions like specialized languages and improved context management [13].
-Furthermore, recent research on LLM in manufacturing delves into language models for specific manufacturing applications.
+The analysis of GPT-4 shows strengths in design and manufacturing knowledge but faces challenges in reasoning and accuracy, with potential solutions including specialized languages and better context management [11].
 Jignasu et al.
-[26] presents a comprehensive evaluation of LLMs for understanding, debugging, and manipulating G-code in 3D printing, testing their capabilities in error detection and correction, geometric transformations, and discussing their potential and limitations in G-code comprehension.
-Similar work has been done by Badini et al.
-[27], which explores using ChatGPT to optimize the G-code generation in additive manufacturing, enhancing efficiency and quality in manufacturing processes.
-These studies prove that LLM can interpret G-code, essential for tasks like debugging, manipulating, and comprehending the code used in manufacturing processes.
-Studies by Fakih et al.
-[28] and Koziolek et al.
-[29] explore the use of LLMs for enhancing programming in industrial control systems, specifically focusing on Programmable Logic Controllers (PLCs).
-Fakih et al.
-developed LLM4PLC, a framework that iteratively refines LLM-generated code through user feedback and verification tools to improve code generation quality.
-Koziolek et al.
-conducted an exploratory study to assess GPT-4’s capability to generate control logic from natural language prompts.
-Several studies have investigated the integration of LLMs into manufacturing, including areas such as CAD/CAM, G-code, and PLCs, aiming to enhance efficiency in manufacturing processes.
+[23] provide a comprehensive evaluation of LLMs for understanding, debugging, and manipulating G-code in 3D printing, testing their capabilities in error detection, correction, and geometric transformations.
+Similar work by Badini et al.
+[24] explores using ChatGPT to optimize the G-code generation in additive manufacturing.
+These studies prove that LLM can interpret G-code, essential for tasks like debugging, manipulating, and comprehending the code.
+Fakih et al.[25] and Koziolek et al.
+[26] focus on enhancing programming in industrial control systems, specifically Programmable Logic Controllers (PLCs), with frameworks like LLM4PLC and studies assessing GPT-4’s capability to generate control logic from natural language prompts.
+Several studies have integrated LLMs into manufacturing, including CAD/CAM, G-code, and PLCs, to enhance efficiency.
 However, these studies have not focused on using LLMs with MAS for autonomous task allocation, adaptability to evolving specifications, and interpretation of context-specific instructions.
-The framework described in Section III focuses on developing autonomous agents that use instructions and specifications in context-specific natural language to dynamically adjust to new product information and coordinate during runtime to complete manufacturing tasks.
-Our proposed framework can be identified as an initial step toward enhancing the CAD/CAM workflow for automatically adapting CAM processes based on changes in CAD specifications, ensuring seamless manufacturing operations.
+The framework described in Section III leverages LLM to enhance multi-agent interactions and decision-making for manufacturing processes.
+This framework develops autonomous agents that use context-specific natural language instructions to dynamically adjust to new product information and coordinate tasks in real-time.
+Our approach is an initial step toward enhancing the CAD/CAM workflow, automatically adapting CAM processes based on changes in CAD specifications, ensuring seamless manufacturing operations.
 III Framework
 Figure 1: Framework for LLM-Enabled Multi-Agent Manufacturing Systems
-The novel architecture presented in Figure 1 is designed based on the architecture of a MAS that includes both PAs and RAs.
+The novel architecture in Figure 1 is designed based on a MAS that includes both Product Agents (PAs) and Resource Agents (RAs).
 A PA is an agent responsible for decision-making related to a product, while an RA serves as a controller for a specific resource, such as a robot or machine [27].
-Figure 1 provides a general overview of an LLM-enabled multi-agent manufacturing system framework, where a user initializes LLM agents with instructions and capabilities.
+Figure 1 provides an overview of an LLM-enabled multi-agent manufacturing system framework, where users initialize agents with instructions and capabilities.
 PAs are assigned specific processes including G-code specifications, while possessing the ability to communicate and access historical data for informed decision-making.
-RAs are responsible for the process execution of tasks, informed by instructions and tool specifications provided by the user.
-LLM is applied for processing and responding to natural language inputs from both PAs and RAs.
+RAs are responsible for the process execution of tasks, informed by instructions and tool specifications.
+LLM is applied for processing natural language inputs from both PAs and RAs.
 For evaluation, the chosen LLM is GPT-4, the latest and most advanced version of GPT developed by OpenAI [8].
@@
 15:  return $Agents$
-The agent initialization process aims to configure a cooperative MAS for operation by defining roles, capabilities, and communication protocols for each agent.
+The agent initialization process configures a MAS by defining roles, capabilities, and communication protocols for each agent.
 This setup ensures agents operate effectively and collaborate towards system goals.
@@
 However, you may call these functions recursively."
-These instructions ensure that each agent operates within the boundaries of its capabilities and knowledge, engages in effective communication, and contributes to the overall objectives of the MAS.
-Following the initial prompts common to all agents in the system, the initialization process involves various parameters, assigning each agent to its specific role.
-These parameters were selected to achieve the desired behaviors of adaptability, real-time response, autonomous decision-making, and efficient coordination within the MAS.
-The process includes the following key parameters essential for the agents:
-  • Functions: Capabilities within the manufacturing process, that range from retrieving manufacturing history to executing manufacturing tasks.
-  • Annotation: Agent’s role and responsibilities, aiding the selection of the appropriate agent among other LLM agents.
+These instructions ensure each agent operates within its capabilities, engages in effective communication, and contributes to the MAS’s objectives.
+Following the initial prompts, the initialization process assigns specific roles to agents using various parameters.
+These parameters are selected to achieve adaptability, real-time response, autonomous decision-making, and efficient coordination within the MAS.
+Key parameters include:
+  • Functions: Capabilities within the manufacturing that range from retrieving history to executing tasks.
+  • Annotation: Agent’s role and responsibilities, which aid in selecting the appropriate LLM agent.
   • Instructions: User instructions offer a sequence of operations, techniques, and standards for agent activities.
-  • Inbox: The agent is initialized with an inbox to receive and manage instructions, updates, and feedback from various agents.
+  • Inbox: The agent receives and manages instructions, updates, and feedback from other agents.
 For PAs, an additional parameter is introduced:
@@
   • Configurations: This parameter outlines the operational settings specific to RAs, such as process times, defect rates, and breakdown probabilities.
-The integration of these parameters into LLM-enabled MAS for manufacturing extends the structured and standardized approaches (e.g.
-FIPA [31] and JADE [32]) to incorporate natural language processing capabilities for more flexible and human-like agent interactions.
-Integrating user-defined instructions and parameters in agent initialization can be further expanded through user studies to ensure applicability to real scenarios and industries.
-Based on the initialization information, the Agent Creation function, cp.
-Algorithm 1, enables the dynamic generation of agents.
-The Agent Creation function provides advanced LLM agents with different capabilities in manufacturing systems, improving operational efficiency and adaptability to changing requirements.
-The algorithm iterates over each entry in the Agent_list, which contains initialization information.
-For every entry, configuration data configuration for the agent is loaded from a JSON, containing details such as the agent’s name, role, capabilities, and any specific functions the agent can perform.
-For each agent specified in the configuration data, an empty list function_list is initially prepared.
-If the ‘functions’ key is present in the agent’s data, the algorithm iterates through each function name listed under this key.
-For each function name, a reference function_reference is obtained by searching within a provided set of agent functions Agent_functions.
-If the function is found, its reference is appended to function_list.
-A new agent instance agent is created with the specified name, agent data, and the list of function references.
-This agent is then added to the Agents list.
-This algorithm facilitates the flexible creation of agents allowing for dynamic adaptation to various tasks and scenarios.
+Integrating parameters into LLM-enabled MAS for manufacturing extends standardized approaches (e.g.
+FIPA [28] and JADE [29]) by incorporating natural language processing for more flexible agent interactions.
+Integrating user-defined instructions and parameters in agent initialization can be further expanded through user studies to ensure industry applicability.
 Algorithm 2 Chat Interaction
@@
 12:  return $function\_response$, $msgs$
+Based on the initialization information, the Agent Creation function, cp.
+Algorithm 1, dynamically generates agents.
+The Agent Creation function provides advanced LLM agents with different capabilities in manufacturing systems for adaptability to changing requirements.
+The algorithm iterates over each entry in the Agent_list, which contains initialization information.
+For every entry, configuration data configuration for the agent is loaded from a JSON, containing details such as the agent’s name, role, capabilities, and any specific functions.
+For each agent, an empty list function_list is prepared.
+If the ‘functions’ key is present, the algorithm iterates through each function name listed.
+For each function name, a reference function_reference is obtained by searching within a provided set of agent functions Agent_functions.
+If the function is found, its reference is appended to function_list.
+A new agent instance agent is created with the specified name, agent data, and the list of function references.
+This agent is then added to the Agents list.
+This algorithm facilitates the flexible creation of agents allowing for adaptation to various tasks.
 III-B Agent Capabilities and LLM Integration
-Through the integration of LLM, agents gain the capability to interpret complex instructions and specifications, facilitating a flexible approach to manufacturing operations.
-The framework employs LLM to enable agents to decode user instructions for manufacturing, integrating a communication mechanism through function information and execution capabilities with the OpenAI Application Programming Interface (API) feature called function calling [33].
-This allows for the automatic execution of user-defined functions, including specific manufacturing operations (e.g., milling, drilling) through descriptions that include operation parameters and required inputs like product name, machine name, process name, and product specifications (e.g.
+Through the integration of LLM, agents gain the capability to interpret complex instructions and specifications, offering a flexible approach to manufacturing operations.
+The framework uses LLM to decode user instructions for manufacturing and leverages the OpenAI Application Programming Interface (API) feature called function calling [30].
+This feature allows for the automatic execution of user-defined functions, including specific manufacturing operations (e.g., milling, drilling) through descriptions that include operation parameters and required inputs like process name and product specifications (e.g.
 G-code).
 The ‘chat’ function, cp.
-Algorithm 2, facilitates dynamic interaction among agents, based on available function information and executing function calls as defined by the manufacturing process requirements, utilizing the appropriate LLM model.
-The outcome, function_response, includes the details like operation type, product, and machine names, along with a message, msgs, to facilitate the subsequent steps in the manufacturing process.
-This approach enables context-aware execution of manufacturing tasks, enhancing the adaptability and efficiency of the manufacturing process.
+Algorithm 2, facilitates interaction among agents, based on available function information and executing function calls as defined by the manufacturing process requirements.
+The outcome, function_response, includes the details like operation type, product, and machine names, along with a message, msgs, to facilitate the subsequent steps in the process.
+This approach enables context-aware execution of tasks, enhancing adaptability and efficiency.
 III-C Operational Workflow
 Figure 2: Operational Workflow of the Framework Using LLM for Task Execution and Coordination
-The workflow, as depicted in Figure 2, shows how user instructions in textual data can be converted into actionable manufacturing operations.
-This approach enables the translation of high-level directives of users into machine-executable actions, enhancing the responsiveness of the manufacturing workflow to changing conditions.
-The workflow begins with the user initializing PAs and RAs within the manufacturing system.
+The workflow, as depicted in Figure 2, shows how user instructions in textual data can be converted into actionable operations.
+This approach enables the translation of high-level directives of users into machine-executable actions, enhancing the responsiveness to changing conditions.
+The workflow begins with the user initializing PAs and RAs.
 Users describe product goals, sequences of operations, and specifications.
@@
 Upon successful operation, the LLM formulates an appropriate response, informing the PA of the process status.
-This feedback updates the LLM-integrated PA with the current process status, determining the completion of the task.
-If additional tasks remain, the cycle repeats; otherwise, the user is notified with the completion of all processes.
+This feedback updates the LLM-integrated PA with the current status, determining the completion of the task.
+If additional tasks remain, the cycle repeats; otherwise, the user is notified of the completion of all processes.
 IV Case Study
-In this case study, we examine the proposed framework and demonstrate how the LLM is integrated into a multi-agent framework to execute the manufacturing process.
+In this case study, we examine the proposed framework and demonstrate how the LLM is integrated into the framework to execute the manufacturing process.
 IV-A Case Study Setup
 Figure 3: Case Study Setup
-A simulated manufacturing environment is equipped with three distinct machines, each specialized for milling, drilling, and threading tasks, as highlighted in Figure 3.
-To evaluate the system’s adaptability to new product information, we structured the case study into two distinct processes: a 2-step process and a more comprehensive 4-step process.
+A simulated manufacturing environment includes three machines specialized for milling, drilling, and threading tasks, as highlighted in Figure 3.
+To evaluate the system’s adaptability, we structured the case study into two distinct processes: a 2-step process and a more comprehensive 4-step process.
 In the 2-step process, a prototype product, referred to as ‘product-1’, is subjected to contouring and drilling operations.
-The 4-step process extends this sequence by incorporating counterboring and threading operations, testing the system’s capability to handle extended manufacturing workflows.
-The G-code specifications for these processes are adapted from [34], as detailed in Figure 4.
+The 4-step process adds counterboring and threading operations, testing the system’s capability to handle extended workflows.
+The G-code specifications are adapted from [31], as detailed in Figure 4.
 Notably, explicit labels for G-code specifications (e.g.
-Step1-Contouring, Step2-Drilling) are not provided to the LLM agent.
-Leveraging LLM’s capability to interpret and assign appropriate G-code based on tool specifications for each resource agent demonstrates the system’s potential to efficiently manage G-code assignments, demonstrating the advanced comprehension capabilities of LLMs in a simulated manufacturing context.
-Each product has an LLM-enabled PA which oversees high-level decision-making for its associated product.
-For instance, ‘product-1’ is instructed for a 2-step process with the following instructions as:
+Step1-Contouring, Step2-Drilling) are not provided to the agent.
+Leveraging LLM’s capability to interpret and assign appropriate G-code based on tool specifications for each RA demonstrates the system’s potential to manage G-code assignments, demonstrating the capabilities of LLMs.
+Each product has an LLM-enabled PA which oversees high-level decision-making for its associated product. For instance, ‘product-1’ is instructed for a 2-step process with the following instructions:
 "Make sure to get contouring to shape the external profile of the product.
 Then, get a drilling operation to create precise holes.
-This operation can precede or follow the contouring.
-Proceed to Exit Buffer completed.
+Proceed to Exit Buffer when completed.
 Your product name is ’product-1’.
 The G-code specifications are as follows: %358 N1 N10 G90 G71 G80 G40..."
-Likewise, each LLM-enabled RA is responsible for operational decisions related to its specific machine.
-‘Milling1’ is tasked with both contouring and counterboring to demonstrate how LLM can adaptively assign operations to the appropriate machine.
-‘Drilling1’ focuses on drilling, while ‘Threading1’ manages threading tasks.
-Each machine is assigned specific tool numbers that are used in G-code specifications to ensure that PAs provide only the relevant G-code instructions to each RA.
-For example, the ‘Milling1’ machine receives the following operational prompt:
+Each LLM-enabled RA handles decisions for its specific machine.
+‘Milling1’ is tasked with both contouring and counterboring to demonstrate how LLM can assign operations adaptively.
+‘Drilling1’ focuses on drilling, while ‘Threading1’ manages threading.
+Machines are assigned tool numbers used in G-code to ensure PAs provide relevant G-code to each RA.
+For example, the ‘Milling1’ machine receives this operational prompt:
 "Perform milling operations as required, including contouring the external shape and counterboring for holes.
 Your machine name is Milling1.
-The tool number for contouring T4.
+The tool number for contouring is T4.
 The tool number for counterboring is T6."
 This experiment investigates the LLM’s capability to identify and allocate correct manufacturing task and their corresponding G-code specifications.
-It aims to enhance the adaptability by showcasing natural language communication between PA and RA, simulating human-like interaction for seamless task execution within the manufacturing environment.
+It aims to enhance the adaptability by showcasing natural language communication between PA and RA, simulating human-like interaction for task execution.
 Figure 4: Case Study G-code Product Specification
@@
 Figure 5: Agent Communication Result for 2-step Process
-Following the setup of our case study, simulations were conducted to evaluate the system performance, as shown in one example, depicted in Figure 5.
-In this example, the user initiates the workflow by instructing the ‘Product-1’ agent and all other RAs, using  Algorithm 1 from Section III-A.
-Following the agent initialization, the system operates according to the steps detailed in the operational workflow, as depicted in Section III-C.
-First, the ‘Product-1’ agent communicates the need for contouring to the ’Milling1’ agent, providing only the contouring section of G-code specifications.
+Following the case study setup, simulations were conducted to evaluate the system, as shown in Figure 5.
+In this example, the user initiates the workflow by instructing the ‘Product-1’ agent and all RAs, using  Algorithm 1 from Section III-A.
+After initialization, the system follows the operational workflow in Section III-C.
+The ‘Product-1’ agent first communicates the need for contouring to the ’Milling1’ agent, providing only the contouring G-code section.
 After the ‘Milling1’ agent completes the task, it informs the ‘Product-1’ agent, indicating it is ready for the next operation.
 The process continues with the ‘Product-1’ agent assigning the drilling operation to the ‘Drilling1’ agent, which updates the ‘Product-1’ agent after completion.
-When all operations are completed, the ‘Product-1’ agent updates the user, indicating that all manufacturing processes for ‘Product-1’ are completed and are now ready to move to the exit buffer.
+When all operations are completed, the ‘Product-1’ agent updates the user, indicating it is ready to move to the exit buffer.
 All agent communications are facilitated by the chat interaction protocol defined in Algorithm 2, as outlined in Section III-B.
-This section details the outcomes from the simulations, highlighting the communications among agents, their understanding and execution of tasks, the system’s accuracy, and the insights gained from these case studies.
 IV-B1 Communication and Comprehension
-Our case study focused on evaluating the agents’ ability to communicate and comprehend specific tasks within the manufacturing process.
-Utilizing natural language processing, agents interacted in both the 2-step and 4-step processes, showcasing their capability for nuanced communication.
-In the 2-step process, the communication between the ’Product-1’ agent and the ‘Milling1’ and ‘Drilling1’ agents showed the system’s efficiency in handling straightforward task sequences and distributing G-code specifications in human-like communication.
-The 4-step process further explored the agents’ capability in managing and understanding complex sequences, requiring accurate G-code allocation where ‘Milling1’ was tasked with both contouring and counterboring.
-In this scenario, the ‘Product-1’ agent coordinated the sequence of contouring, drilling, counterboring, and threading, showing the agents’ ability to interpret and implement longer instructions.
+Our case study evaluated the agents’ ability to communicate and comprehend specific tasks.
+Using natural language processing, agents interacted in both 2-step and 4-step processes.
+In the 2-step process, the communication between the ’Product-1’ agent and the ‘Milling1’ and ‘Drilling1’ agents showed the system’s efficiency in handling straightforward task sequences and distributing G-code specifications.
+The 4-step process tested the agents’ ability to manage complex sequences, with ‘Milling1’ handling both contouring and counterboring.
+In this scenario, the ‘Product-1’ agent coordinated contouring, drilling, counterboring, and threading, showing the agents’ ability to interpret complex instructions.
 IV-B2 Accuracy and Performance
-The data presented in Table I reveals that GPT-4 achieved success rates of 100% for 2-step and 86% for 4-step manufacturing processes across 50 trials for each.
-This shows a decline in performance as task complexity increases, with the 4-step process experiencing a 14% overall error rate.
-The detailed error breakdown in  Table II reveals that within the overall 14% error rate for the 4-step process, incorrect function calls and process inexecution have 43% of the errors, while inaccurate G-code allocation accounts for 14% of the total errors.
-It is observed that the increase in prompt length correlates with a rise in errors, highlighting the need for refined model training and improved prompt engineering.
-Addressing these issues aims to enhance task execution capabilities and effective integration of LLMs into MASs for manufacturing, improving system adaptability and operational efficiency even in unexpected scenarios.
+The data in Table I shows GPT-4 achieved 100% success in 2-step and 86% in 4-step processes across 50 trials each, with performance declining as task complexity increased.
+The 4-step process had a 14% overall error rate.
+The detailed error breakdown in Table II shows that in the 4-step process, 43% of errors were due to incorrect function calls and process failures, and 14% were due to inaccurate G-code allocation.
+The increase in prompt length correlates with a rise in errors, highlighting the need for model training and improved prompt engineering.
+Addressing these issues will enhance task execution capabilities in MASs for manufacturing, improving system adaptability and efficiency even in unexpected scenarios.
 IV-B3 Insights from Case Studies
 The case study provided insights into the capabilities of LLM-enabled MASs in manufacturing.
-The current framework incorporates LLM to transition towards a system capable of understanding the manufacturing tasks (Cn1), interpreting and adapting to new product specifications (Cn2), and agent communication via natural language instructions (Cn3).
-This shift enhances system flexibility and expands its ability to manage various tasks without relying on predetermined models or rules.
-However, accuracy issues with GPT-4 highlight the need for model improvement and adaptation to manufacturing needs.
-Difficulties with incorrect argument passing during function calls point to the necessity for better prompt engineering, error management, and validation within the system.
-These results guide future research toward enhancing LLM integration in manufacturing, developing advanced agent communication protocols, and increasing system robustness against errors.
+The framework uses LLM to understand manufacturing tasks (Cn1), adapt to new product specifications (Cn2), and facilitate agent communication via natural language (Cn3).
+This enhances system flexibility and task management without relying on predetermined rules.
+However, accuracy issues with GPT-4 highlight the need for model improvement to manufacturing needs.
+Difficulties with incorrect function calls point to the necessity for better error management and validation.
+These results guide future research toward enhancing LLM integration, developing advanced agent communication protocols, and increasing system robustness.
 TABLE I: Success Rates for Case Studies with GPT-4
@@
 V-A Performance
-A significant challenge within the use of LLMs in manufacturing processes is achieving the level of performance required for high-quality production.
-Manufacturing processes demand precision to ensure product consistency, safety, and reliability.
-However, LLMs, while effective in processing natural language, occasionally shows limitations in delivering the precision necessary for these applications, as indicated in Section IV-B2.
-This issue raises concerns about the feasibility of relying on LLMs for tasks where errors can lead to other consequences, such as machine breakdown and low product quality.
+A significant challenge in manufacturing using LLMs is achieving the performance required for high-quality production.
+While effective in processing natural language, LLMs sometimes lack precision, as indicated in Section IV-B2.
+This issue raises concerns about the feasibility of relying on LLMs for tasks where errors can cause machine breakdown and low product quality.
+The proposed framework can enhance performance by applying fine-tuned LLMs on domain-specific data [32].
 V-B Scalability
-Scalability is also a challenge when integrating LLMs into manufacturing systems.
-As the volume of data and the complexity of tasks grow, there is a risk that LLMs may not maintain their initial levels of accuracy and efficiency.
-This limitation is particularly noticeable in situations where extensive historical data and detailed prompts are essential for task execution, as exemplified in the 4-step process in Section IV.
+As data volume and task complexity grow, LLMs may struggle to maintain accuracy.
+This limitation is particularly noticeable in situations where extensive historical data and detailed prompts are essential, as exemplified in the 4-step process in Section IV.
 The potential decrease in accuracy with larger datasets poses a challenge to scaling LLM-enabled multi-agent control across more complex manufacturing operations.
+A knowledge base for multi-agent manufacturing systems [33] should be integrated for practical industrial applications to provide LLMs with accurate runtime data and to enhance scalability.
 V-C Reliability
-The integration of LLMs into manufacturing settings highlights the importance of reliability.
-LLMs generating hallucinated responses or making false decisions raises significant safety concerns, as shown in Section IV.
-The LLM in the case study allocated inaccurate G-code or executed the incorrect function calls.
+Reliability is key for integrating this method into real manufacturing scenarios.
+LLMs generating hallucinated responses or making false decisions raises safety concerns, as shown in Section IV.
+The LLM allocated inaccurate G-code or executed the incorrect function calls.
 Moreover, the reasoning provided by LLMs may require validation in practical manufacturing scenarios even if it seems logical.
@@
 This approach ensures LLMs’ outputs and reasoning are consistent and allow immediate adjustments with manufacturing standards and safety protocols.
-For instance, the study by Joublin et al.
-[35] presents a system architecture for robotics that utilizes LLMs for planning tasks, focusing on a strategy for replanning in response to errors via feedback.
-This method enables robots to adjust to failures and changes, enhancing reliability.
-Thus, feedback loops in LLM systems ensure safety and reliability in complex manufacturing environments.
+Joublin et al.
+[34] demonstrate a system for robotics that uses LLMs for task planning and replanning in response to errors, highlighting the importance of feedback loops for safety and reliability in complex manufacturing environments.
 VI Conclusion and Future Work
-In our research, we explored the integration of LLM within MAS to enhance manufacturing processes.
-Our research demonstrates the effect of LLM in manufacturing, where they serve to dynamically match tasks with optimal process parameters, improving task allocation efficiency and operational adaptability (Cn1).
-Moreover, LLM enables agents to adapt to evolving product specifications (e.g.
-G-code), meeting the manufacturing requirements (Cn2).
-Additionally, through advanced natural language processing, LLM enables agents to effectively understand and execute complex, context-specific instructions from humans and other agents enhancing the agent communication and collaboration within the manufacturing process (Cn3).
-However, challenges have been identified, such as task execution performance, scalability, and reliability, highlighting the need for a verification process and monitoring capabilities when making decisions for manufacturing operations.
-Furthermore, the diversity in user instructions based on different skill levels and manufacturing requirements affects task interpretation and execution, which requires a strategy to dynamically adapt to these changes when integrating LLM to ensure reliability.
-Future work will focus on developing communication protocols incorporating LLM for agents that can accommodate a wide variety of manufacturing instructions and requirements, influenced by varying proficiency levels of human operators.
-This approach seeks to showcase the capability of agents to learn and adapt from historical data, enhancing their responsiveness to rapid changes in modern manufacturing industries.
+In our research, we explored integrating LLM within MAS to enhance manufacturing processes.
+Our research demonstrates that LLM improves task allocation efficiency and operational adaptability (Cn1), enables agents to adapt to evolving product specifications (e.g., G-code) (Cn2), and enhances agent communication and collaboration through advanced natural language processing (Cn3).
+However, challenges like task execution performance, scalability, and reliability highlight the need for verification and monitoring.
+Additionally, the diversity in user instructions requires a strategy to ensure reliable task interpretation and execution.
+Future work will focus on developing LLM-based communication protocols for agents to handle diverse manufacturing instructions and operator proficiency levels.
+This approach seeks to showcase the agents’ capability to learn and adapt from historical data, enhancing their responsiveness to rapid industry changes.
 Additionally, efforts will be directed towards enhancing the accuracy and scalability of LLM.
-While this paper initially focused on widely used manufacturing processes (e.g.
-drilling, milling), future investigations will extend to include less commonly used manufacturing processes with larger datasets to evaluate the effectiveness of our framework.
-The goal is to assess if integrating additional processes results in less effective outcomes or if they can also be improved by applying these methodologies, ensuring efficient processing of larger datasets and diverse manufacturing processes without impacting performance.
+While this paper focused on common manufacturing processes (e.g.
+drilling, milling), future research will include less common processes with larger datasets to evaluate the framework’s effectiveness.
+The goal is to ensure the efficient processing of diverse manufacturing tasks without impacting performance.
 Acknowledgment